find a quadratic equation with roots (4+i) and (4-i)
Answer by stanbon(48568)

find a quadratic equation with roots (4+i) and (4-i)
-----
f(x) = (x-(4+i))(x-(4-i))
---
f(x) = ((x-4)-i)((x-4)+i)
---
f(x) = (x-4)^2 - i^2
---
f(x) = (x^2-8x+16 + 1
f(x) = x^2 - 8x + 17
============================
Cheers,
Stan H.